# Winner-takes-all for Multivariate Probabilistic Time Series Forecasting

### v1

@@ 3 Problem setup and notations
Let $(x_{t:T})\in\mathcal{X}^{T-t+1}$ represent a multivariate time series on $\mathcal{X}=\mathbb{R}^{D}$ over time indexes $[\![t,T]\!]$, where $t\in[\![1,T]\!]$. We aim to learn the conditional law
@@
of future values $x_{t_{0}:T}$ (over the interval $[\![t_{0},T]\!]$) given past observations $x_{1:t_{0}-1}$ and covariates $c_{1:T}$, the latter being omitted in the following for conciseness. Our focus is on scenarios where the conditional distribution may exhibit multiple modes (multi-modality), motivating a richer representation than a single-mean regressor.
@@ 4.3 Taking advantage of Winner-takes-all variants
~$\tilde{\mathcal{L}}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{k})\triangleq\mathbb{E}_{x_{1:T}}\left[\sum_{k=1}^{K}q_{k}\mathcal{L}^{k}_{\theta}(x_{1:t_{0}-1},x_{t_{0}:T})\right]\;,$ (6)
@@ 5.1 TimeMCL is a stationary conditional functional quantizer
[…]
@@ 6.2 Results
Elec. 14041 $\pm$ 877 13407 $\pm$ 2921 9558 $\pm$ 1024 9799 $\pm$ 660 133640 $\pm$ 2341 14787 $\pm$ 674 11287 $\pm$ 730 11493 $\pm$ 1326
Exch. 0.051 $\pm$ 0.005 0.062 $\pm$ 0.015 0.031 $\pm$ 0.002 0.035 $\pm$ 0.005 0.061 $\pm$ 0.001 0.051 $\pm$ 0.008 0.033 $\pm$ 0.004 0.041 $\pm$ 0.003
Solar 641.32 $\pm$ 3.57 376.98 $\pm$ 21.85 362.21 $\pm$ 39.94 360.78 $\pm$ 34.57 746.12 $\pm$ 24.91 368.93 $\pm$ 18.0 280.66 $\pm$ 22.23 286.3 $\pm$ 21.59
Traffic 2.64 $\pm$ 0.01 1.26 $\pm$ 0.01 0.85 $\pm$ 0.03 0.77 $\pm$ 0.01 2.13 $\pm$ 0.05 1.19 $\pm$ 0.02 0.68 $\pm$ 0.01 0.72 $\pm$ 0.05
Taxi 583.52 $\pm$ 0.41 277.96 $\pm$ 13.0 246.09 $\pm$ 11.32 208.32 $\pm$ 2.76 407.63 $\pm$ 15.89 268.18 $\pm$ 12.18 186.99 $\pm$ 6.15 228.18 $\pm$ 21.3
Wiki. 715150 $\pm$ 4742 515285 $\pm$ 16000 254109 $\pm$ 4322 260927 $\pm$ 1185 366286 $\pm$ 8924 382819 $\pm$ 3508 258256 $\pm$ 3253 260025 $\pm$ 2649
Table 2: Computational cost at inference of neural-based methods for $K=16$ hypotheses on Exchange. FLOPs are computed with a single batch of size $1$. Run-times are averaged over $15$ random seeds.
@@
Comparing TimeMCL with the baselines on standard metrics. Table 7 and 8 provides performance on CRPS and RMSE, respectively. We see that, except for CRPS and RMSE on the Solar dataset and RMSE on exchange (for which Tactis is better), TimeGrad outperforms the other baselines. We also observe that TimeMCL is competitive despite optimizing a completely different training objective, at a fraction of TimeGrad’s and Tactis’s computational cost.
@@
Elec. 276601 $\pm$ 19341 371934 $\pm$ 16589 433494 $\pm$ 23503 220372 $\pm$ 29409 245904 $\pm$ 26410
Exch. 0.225 $\pm$ 0.0176 0.606 $\pm$ 0.016 1.156 $\pm$ 0.155 0.031 $\pm$ 0.0065 0.042 $\pm$ 0.0123
Solar 4249 $\pm$ 240 6505 $\pm$ 1897 3637 $\pm$ 354 3391 $\pm$ 820 2195 $\pm$ 257
Traffic 10.89 $\pm$ 1.079 12.9 $\pm$ 1.391 17.644 $\pm$ 0.234 5.766 $\pm$ 0.165 5.715 $\pm$ 0.245
Taxi 5262.07 $\pm$ 592.83 3728.29 $\pm$ 6.39 4104.29 $\pm$ 361.39 712.56 $\pm$ 291.79 703.29 $\pm$ 171.0
Wiki. 2650865 $\pm$ 422885 2518198 $\pm$ 233471 12729658 $\pm$ 367297 9530 $\pm$ 9256 271611 $\pm$ 311180
@@ 6.3 Effect of the number of hypotheses
Table 4: Effect of the number of hypotheses on the Distortion Risk for Solar. Results are averaged over three training seeds. TimeMCL(R.) and TimeMCL(A.) corresponds to the relaxed and annealed variants. See Table 9 for the full results.
@@ Acknowledgements
This work was funded by the French Association for Technological Research (ANRT CIFRE contract 2022-1854), and granted access to the HPC resources of IDRIS (allocation 2024-AD011014345) by GENCI. We are grateful to the reviewers for their insightful comments.
@@ Appendix A Theoretical results
$-\int_{\mathcal{X}_{k}(x_{1:t_{0}-1})}\log\bigl{(}\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}\,p(x_{t_{0}:T}\mid x_{1:t_{0}-1})\,\mathrm{d}x_{t_{0}:T}~{}-\int_{\mathcal{X}^{t_{0}:T}\setminus\mathcal{X}_{k}(x_{1:t_{0}-1})}\log\bigl{(}1-\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}\,p(x_{t_{0}:T}\mid x_{1:t_{0}-1})\,\mathrm{d}x_{t_{0}:T}.$
@@
$-\,m_{k}(x_{1:t_{0}-1})\,\log\bigl{(}\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}~{}-\,(1-m_{k}(x_{1:t_{0}-1}))\,\log\bigl{(}1-\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}\;.$
@@ Appendix B Experimental details on the synthetic datasets
In order to conduct interpretable experiments and generate synthetic data, we employed Brownian motion. A Brownian motion $\{W_{t}\}_{t\geq 0}$ is a Gaussian process characterized by a mean $\mathbb{E}[W_{t}]=0$ and a covariance function $\mathrm{Cov}(W_{t},W_{s})=\min(t,s)$ for $t,s\geq 0$. This process possesses several properties of interest (Karatzas et al., 1998). In particular, the Markov property, which ensures that the process $\{W_{s+t}-W_{s}\}_{t\geq 0}$ is itself a Brownian motion starting at 00, independent of the process $\{W_{t}:0\leq t\leq s\}$. This implies that to condition a Brownian motion on its trajectory up to time $t$, it suffices to condition on its value $W_{t}$. The process beyond time $t$ behaves as if it were a new Brownian motion originating at $W_{t}$. In conditional quantization, we aim to quantize Brownian motion at a given prediction horizon, conditioned on its value at time $t$. To compute the target quantization, we followed the methodology detailed in (Pages & Printems, 2009), leveraging the Karhunen-Loève decomposition (Arcozzi et al., 2015) of the Brownian motion on $t\in[0,1]$:
@@ C.2.2 CRPS (Continuous Ranked Probability Score)
$\mathscr{L}_{q}\bigl{(}x_{t_{0}:T}^{i},\hat{x}_{t_{0}:T}^{i}\bigr{)}$ $\triangleq 2\sum_{t=t_{0}}^{T}\left\lvert\mathcal{A}({x}_{t}^{i})-Q_{X_{t}^{i}}(q)\right\rvert\!\left(\mathds{1}(\mathcal{A}({x}_{t}^{i})\leq Q_{X_{t}^{i}}(q))-q\right),\quad\mathscr{T}(x_{t_{0}:T})\triangleq\frac{1}{N}\sum_{i=1}^{N}\sum_{t=t_{0}}^{T}\left\lvert\mathcal{A}({x}_{t}^{i})\right\rvert,$
@@ C.3 Experimental details
Elec. 15132 $\pm$ 1662 14049 $\pm$ 3492 9922 $\pm$ 938 10165 $\pm$ 750 143638 $\pm$ 5048 15298 $\pm$ 760 12156 $\pm$ 415 11371 $\pm$ 951
Exch. 0.057 $\pm$ 0.002 0.066 $\pm$ 0.012 0.032 $\pm$ 0.002 0.038 $\pm$ 0.006 0.062 $\pm$ 0.001 0.052 $\pm$ 0.008 0.042 $\pm$ 0.007 0.042 $\pm$ 0.006
Solar 654.48 $\pm$ 3.45 385.09 $\pm$ 16.33 370.1 $\pm$ 37.09 371.17 $\pm$ 37.34 768.19 $\pm$ 22.89 379.16 $\pm$ 24.42 311.72 $\pm$ 22.4 334.02 $\pm$ 26.74
Traffic 2.65 $\pm$ 0.0 1.29 $\pm$ 0.03 0.86 $\pm$ 0.03 0.77 $\pm$ 0.01 2.18 $\pm$ 0.06 1.22 $\pm$ 0.03 0.69 $\pm$ 0.02 0.71 $\pm$ 0.02
Taxi 589.15 $\pm$ 0.75 283.73 $\pm$ 14.62 247.4 $\pm$ 11.41 209.94 $\pm$ 2.85 425.09 $\pm$ 16.74 273.27 $\pm$ 11.43 188.78 $\pm$ 5.53 254.46 $\pm$ 22.83
Wiki. 722166 $\pm$ 5248 529398 $\pm$ 14698 254772 $\pm$ 4581 261326 $\pm$ 1023 368717 $\pm$ 6238 382839 $\pm$ 3959 264596 $\pm$ 16414 257624 $\pm$ 2931
@@ C.3.1 Architectures, training, and inference
Elec. 25771 $\pm$ 1117 20551 $\pm$ 11019 15897 $\pm$ 2848 12060 $\pm$ 905 147404 $\pm$ 1653 23590 $\pm$ 2787 17963 $\pm$ 2619 18208 $\pm$ 3070
Exch. 0.089 $\pm$ 0.01 0.147 $\pm$ 0.043 0.071 $\pm$ 0.009 0.078 $\pm$ 0.026 0.085 $\pm$ 0.004 0.093 $\pm$ 0.018 0.089 $\pm$ 0.012 0.086 $\pm$ 0.014
Solar 3661.86 $\pm$ 28.87 4112.38 $\pm$ 239.2 3317.51 $\pm$ 671.32 3193.21 $\pm$ 324.15 7181.85 $\pm$ 127.04 3793.0 $\pm$ 62.22 3100.51 $\pm$ 321.95 3622.0 $\pm$ 573.02
Traffic 15.24 $\pm$ 0.09 27.63 $\pm$ 0.27 12.17 $\pm$ 0.71 3.92 $\pm$ 0.47 39.73 $\pm$ 1.3 26.65 $\pm$ 1.19 5.46 $\pm$ 0.65 5.72 $\pm$ 0.69
Taxi 9779.81 $\pm$ 13.42 4223.77 $\pm$ 565.52 2826.4 $\pm$ 184.3 2321.4 $\pm$ 113.16 7395.54 $\pm$ 273.7 4428.98 $\pm$ 595.0 6789.85 $\pm$ 3423.68 8903.44 $\pm$ 2367.2
Wiki. 765663 $\pm$ 39186 1152828 $\pm$ 171868 584384 $\pm$ 113710 501638 $\pm$ 59512 3562564 $\pm$ 1802493 963681 $\pm$ 142202 835145 $\pm$ 138204 3623632 $\pm$ 2956320
@@
Elec. 0.0601 $\pm$ 0.0045 0.0433 $\pm$ 0.0214 0.0341 $\pm$ 0.0056 0.0241 $\pm$ 0.001 0.424 $\pm$ 0.0099 0.0503 $\pm$ 0.0043 0.0473 $\pm$ 0.0066 0.047 $\pm$ 0.0091
Exch. 0.0079 $\pm$ 0.001 0.0138 $\pm$ 0.0039 0.0077 $\pm$ 0.0013 0.0075 $\pm$ 0.0028 0.008 $\pm$ 0.0003 0.0088 $\pm$ 0.0017 0.009 $\pm$ 0.0015 0.0104 $\pm$ 0.0017
Solar 0.509 $\pm$ 0.0041 0.4524 $\pm$ 0.0256 0.3785 $\pm$ 0.0648 0.3682 $\pm$ 0.0442 1.0235 $\pm$ 0.0286 0.4352 $\pm$ 0.0139 0.3518 $\pm$ 0.0371 0.3615 $\pm$ 0.0365
Traffic 0.2007 $\pm$ 0.0014 0.3963 $\pm$ 0.0065 0.143 $\pm$ 0.0126 0.0488 $\pm$ 0.005 0.5295 $\pm$ 0.0216 0.3812 $\pm$ 0.0255 0.0664 $\pm$ 0.0089 0.0657 $\pm$ 0.0091
Taxi 0.8948 $\pm$ 0.0017 0.2872 $\pm$ 0.0467 0.1975 $\pm$ 0.0165 0.1417 $\pm$ 0.0061 0.5262 $\pm$ 0.0245 0.3113 $\pm$ 0.0533 0.4429 $\pm$ 0.1916 0.5965 $\pm$ 0.2101
Wiki. 0.0789 $\pm$ 0.0041 0.1309 $\pm$ 0.0316 0.0682 $\pm$ 0.0167 0.0562 $\pm$ 0.0084 0.5193 $\pm$ 0.2987 0.0964 $\pm$ 0.0079 0.0896 $\pm$ 0.0233 0.1325 $\pm$ 0.026
Optimization Training is conducted using the Adam optimizer with a learning rate of $10^{-3}$, and following an ‘LR on plateau’ scheduler (See Pytorch documentation), a weight decay of $10^{-8}$ and $200$ training epochs. Additionally, a separate validation split is used with a temporal size equal to $10$ times the prediction length. When training TimeMCL, the WTA loss (and its variants) was divided by the prediction length $T-t_{0}+1$. All models are trained and evaluated using Pytorch float32 precision.

### v2

@@ 3 Problem setup and notations
Let $(x_{t:T})\in\mathcal{X}^{T-t+1}$ represent a multivariate time series on $\mathcal{X}=\mathbb{R}^{D}$ over time indexes [[t,T]][\![t,T]\!], where $t\in[\![1,T]\!]$. We aim to learn the conditional law
@@
of future values $x_{t_{0}:T}$ (over the interval $t_{0},T]\!]$) given past observations $x_{1:t_{0}-1}$ and covariates $c_{1:T}$, the latter being omitted in the following for conciseness. Our focus is on scenarios where the conditional distribution may exhibit multiple modes (multi-modality), motivating a richer representation than a single-mean regressor.
@@ 4.3 Taking advantage of Winner-takes-all variants
$\tilde{\mathcal{L}}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{k})\triangleq\mathbb{E}_{x_{1:T}}\left[\sum_{k=1}^{K}q_{k}\mathcal{L}^{k}_{\theta}(x_{1:t_{0}-1},x_{t_{0}:T})\right]\;,$ (6)
@@ 5.1 TimeMCL is a stationary conditional functional quantizer
For simplicity and without loss of generality, let us temporarily assume that $\beta=0$, i.e., only the WTA Loss $\mathcal{L}^{\mathrm{WTA}}$ is optimized. When predicting the future of a time series given its context $x_{1:t_{0}-1}$, we are effectively observing, during supervision, a (partial) path realization $x_{t_{0}:T}$ of an underlying stochastic process. Following standard functional data analysis (Bosq, 2000; Ramsay & Silverman, 2005), we assume that each time-series trajectory $x_{t_{0}:T}$ belongs to a Banach space $(E,\|\cdot\|)$. For concreteness, one may consider $E=L^{2}([t_{0},T])$ endowed with the usual $L^{2}$ norm, though any separable Banach space is admissible for the theoretical arguments to hold. In this space, the distance $d(\cdot,\cdot)$ induced by $\|\cdot\|$ enables us to define the Voronoi tessellation over future paths $x_{t_{0}:T}$ as described below.
@@ 6.2 Results
Elec. 14041 $\pm$ 877 13519 $\pm$ 2789 11616 $\pm$ 2160 9872 $\pm$ 668 133107 $\pm$ 2870 14836 $\pm$ 596 11604 $\pm$ 1042 11611 $\pm$ 1206
Exch. 0.051 $\pm$ 0.005 0.062 $\pm$ 0.015 0.030 $\pm$ 0.002 0.038 $\pm$ 0.005 0.061 $\pm$ 0.001 0.051 $\pm$ 0.007 0.035 $\pm$ 0.005 0.044 $\pm$ 0.005
Solar 641.32 $\pm$ 3.57 374.69 $\pm$ 20.97 358.01 $\pm$ 38.47 360.6 $\pm$ 34.79 748.68 $\pm$ 18.92 371.14 $\pm$ 18.2 280.03 $\pm$ 15.22 305.5 $\pm$ 4.05
Traffic 2.64 $\pm$ 0.01 1.26 $\pm$ 0.02 0.84 $\pm$ 0.04 0.78 $\pm$ 0.05 2.12 $\pm$ 0.08 1.21 $\pm$ 0.04 0.68 $\pm$ 0.01 0.72 $\pm$ 0.05
Taxi 583.52 $\pm$ 0.41 278.22 $\pm$ 13.8 243.63 $\pm$ 9.1 209.64 $\pm$ 3.14 407.4 $\pm$ 14.09 268.65 $\pm$ 10.69 187.81 $\pm$ 6.17 229.26 $\pm$ 22.06
Wiki 715150 $\pm$ 4742 515285 $\pm$ 16000 254675 $\pm$ 3870 261361 $\pm$ 1028 368087 $\pm$ 8159 382554 $\pm$ 4149 261950 $\pm$ 9504 267624 $\pm$ 8433
Table 2: Computational cost at inference of neural-based methods for $K=16$ hypotheses on Exchange. FLOPs are computed with a single batch of size 11. Run-times are averaged over $15$ random seeds.
@@
Comparing TimeMCL with the baselines on standard metrics. Table 7 and 8 provides performance on CRPS and RMSE, respectively. We see that, except for CRPS and RMSE on exchange (for which Tactis is better), TimeGrad outperforms the other baselines. We also observe that TimeMCL is competitive despite optimizing a completely different training objective, at a fraction of TimeGrad’s and Tactis’s computational cost.
@@
Elec. 284232 $\pm$ 23269 372443 $\pm$ 18499 434097 $\pm$ 34623 220375 $\pm$ 33961 245908 $\pm$ 30505
Exch. 0.237 $\pm$ 0.0355 0.606 $\pm$ 0.0208 1.104 $\pm$ 0.1582 0.031 $\pm$ 0.0075 0.042 $\pm$ 0.0141
Solar 4421 $\pm$ 828 5383 $\pm$ 2173 3653 $\pm$ 488 3391 $\pm$ 947 2195 $\pm$ 297
Traffic 11.063 $\pm$ 0.874 12.991 $\pm$ 1.661 18.114 $\pm$ 0.512 5.766 $\pm$ 0.19 5.714 $\pm$ 0.283
Taxi 5452.77 $\pm$ 324.07 4232.16 $\pm$ 713.9 4147.2 $\pm$ 362.87 709.86 $\pm$ 332.07 701.61 $\pm$ 199.35
Wiki 2634547 $\pm$ 597323 2458503 $\pm$ 151593 12410909 $\pm$ 319517 9532 $\pm$ 10690 271611 $\pm$ 359320
@@ 6.3 Effect of the number of hypotheses
Table 4: Effect of the number of hypotheses on the Distortion Risk for Solar. Results are averaged over four training seeds. TimeMCL(R.) and TimeMCL(A.) corresponds to the relaxed and annealed variants. See Table 9 for the full results.
@@ Acknowledgements
This work was funded by the French Association for Technological Research (ANRT CIFRE contract 2022-1854) and the LISTEN Laboratory of Télécom Paris. It also benefited from access to the HPC resources of IDRIS (allocation 2024-AD011014345) by GENCI. We are grateful to the reviewers for their insightful comments.
@@ Appendix A Theoretical results
$\int_{\mathcal{X}_{k}(x_{1:t_{0}-1})}\log\bigl{(}\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}\,p(x_{t_{0}:T}\mid x_{1:t_{0}-1})\,\mathrm{d}x_{t_{0}:T}~-\int_{\mathcal{X}^{t_{0}:T}\setminus\mathcal{X}_{k}(x_{1:t_{0}-1})}\log\bigl{(}1-\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}\,p(x_{t_{0}:T}\mid x_{1:t_{0}-1})\,\mathrm{d}x_{t_{0}:T}.$
@@
$m_{k}(x_{1:t_{0}-1})\,\log\bigl{(}\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}~-\,(1-m_{k}(x_{1:t_{0}-1}))\,\log\bigl{(}1-\Gamma_{\theta}^{k}(x_{1:t_{0}-1})\bigr{)}\;.$
@@ Appendix B Experimental details on the synthetic datasets
In order to conduct interpretable experiments and generate synthetic data, we employed Brownian motion. A Brownian motion $\{W_{t}\}_{t\geq 0}$ is a Gaussian process characterized by a mean $\mathbb{E}[W_{t}]=0$ and a covariance function $\mathrm{Cov}(W_{t},W_{s})=\min(t,s)$ for $t,s\geq 0$. This process possesses several properties of interest (Karatzas et al., 1998). In particular, the Markov property, which ensures that the process $\{W_{s+t}-W_{s}\}_{t\geq 0}$ is itself a Brownian motion starting at 0, independent of the process $\{W_{t}:0\leq t\leq s\}$. This implies that to condition a Brownian motion on its trajectory up to time $t$, it suffices to condition on its value $W_{t}$. The process beyond time $t$ behaves as if it were a new Brownian motion originating at $W_{t}$. In conditional quantization, we aim to quantize Brownian motion at a given prediction horizon, conditioned on its value at time $t$. To compute the target quantization, we followed the methodology detailed in (Pages & Printems, 2009), leveraging the Karhunen-Loève decomposition (Arcozzi et al., 2015) of the Brownian motion on $t\in[0,1]$:
@@ C.2.2 CRPS (Continuous Ranked Probability Score)
$\mathscr{L}_{q}\bigl{(}x_{t_{0}:T}^{i},\hat{x}_{t_{0}:T}^{i}\bigr{)}$ $\triangleq 2\sum_{t=t_{0}}^{T}\left\lvert\mathcal{A}({x}_{t}^{i})-Q_{X_{t}^{i}}(q)\right\rvert\!\left(\mathds{1}(\mathcal{A}({x}_{t}^{i})\leq Q_{X_{t}^{i}}(q))-q\right),\quad\mathscr{T}(x_{t_{0}:T})\triangleq\sum_{i=1}^{N}\sum_{t=t_{0}}^{T}\left\lvert\mathcal{A}({x}_{t}^{i})\right\rvert$,
@@ C.3 Experimental details
Elec. 15132 $\pm$ 1662 14160 $\pm$ 3370 12557 $\pm$ 2405 10172 $\pm$ 680 144832 $\pm$ 3081 15822 $\pm$ 531 12155 $\pm$ 414 11509 $\pm$ 703
Exch. 0.057 $\pm$ 0.002 0.066 $\pm$ 0.012 0.031 $\pm$ 0.002 0.041 $\pm$ 0.006 0.062 $\pm$ 0.001 0.051 $\pm$ 0.007 0.039 $\pm$ 0.001 0.042 $\pm$ 0.005
Solar 654.48 $\pm$ 3.45 387.15 $\pm$ 18.57 367.54 $\pm$ 35.4 371.97 $\pm$ 33.31 770.48 $\pm$ 19.91 379.93 $\pm$ 22.38 305.63 $\pm$ 28.24 347.82 $\pm$ 28.82
Traffic 2.65 $\pm$ 0.0 1.28 $\pm$ 0.02 0.85 $\pm$ 0.04 0.79 $\pm$ 0.05 2.14 $\pm$ 0.06 1.23 $\pm$ 0.04 0.69 $\pm$ 0.02 0.71 $\pm$ 0.01
Taxi 589.15 $\pm$ 0.75 284.25 $\pm$ 13.94 245.09 $\pm$ 8.79 211.71 $\pm$ 4.02 429.01 $\pm$ 14.88 272.84 $\pm$ 12.14 187.04 $\pm$ 2.45 256.16 $\pm$ 24.95
Wiki 722166 $\pm$ 5248 528546 $\pm$ 11657 255328 $\pm$ 3386 261463 $\pm$ 1047 371035 $\pm$ 6578 388511 $\pm$ 2472 271878 $\pm$ 17358 258468 $\pm$ 3108
@@ C.3.1 Architectures, training, and inference
Elec. 25771 $\pm$ 1117 20559 $\pm$ 11011 19342 $\pm$ 5496 11597 $\pm$ 931 145977 $\pm$ 3564 23373 $\pm$ 2614 18571 $\pm$ 3063 19154 $\pm$ 2657
Exch. 0.089 $\pm$ 0.01 0.147 $\pm$ 0.043 0.062 $\pm$ 0.007 0.077 $\pm$ 0.027 0.086 $\pm$ 0.002 0.095 $\pm$ 0.021 0.084 $\pm$ 0.002 0.103 $\pm$ 0.021
Solar 3661.86 $\pm$ 28.87 4121.74 $\pm$ 243.06 3531.02 $\pm$ 439.59 3244.92 $\pm$ 247.16 7199.09 $\pm$ 107.71 3801.72 $\pm$ 122.09 3639.23 $\pm$ 606.34 3896.48 $\pm$ 468.71
Traffic 15.24 $\pm$ 0.09 27.57 $\pm$ 0.22 11.54 $\pm$ 1.29 4.44 $\pm$ 1.07 39.15 $\pm$ 1.48 26.77 $\pm$ 1.29 5.48 $\pm$ 0.71 5.73 $\pm$ 0.7
Taxi 9779.81 $\pm$ 13.42 4223.88 $\pm$ 561.11 2662.52 $\pm$ 58.79 2357.85 $\pm$ 133.05 7323.32 $\pm$ 245.35 4436.39 $\pm$ 581.06 6855.32 $\pm$ 3456.01 8549.15 $\pm$ 2691.98
Wiki 765663 $\pm$ 39186 1152828 $\pm$ 171868 564176 $\pm$ 115593 505869 $\pm$ 60432 4469007 $\pm$ 1015212 864607 $\pm$ 105223 1210500 $\pm$ 412274 1264339 $\pm$ 166347
@@
Elec. 0.0601 $\pm$ 0.0045 0.0434 $\pm$ 0.0214 0.0409 $\pm$ 0.0105 0.0235 $\pm$ 0.0018 0.4234 $\pm$ 0.0101 0.0503 $\pm$ 0.0038 0.0488 $\pm$ 0.0074 0.0481 $\pm$ 0.0086
Exch. 0.0079 $\pm$ 0.001 0.0138 $\pm$ 0.0039 0.0066 $\pm$ 0.0014 0.0074 $\pm$ 0.0029 0.0081 $\pm$ 0.0002 0.0091 $\pm$ 0.0019 0.0103 $\pm$ 0.0002 0.0131 $\pm$ 0.0031
Solar 0.509 $\pm$ 0.0041 0.4528 $\pm$ 0.0271 0.3919 $\pm$ 0.0378 0.3789 $\pm$ 0.0356 1.028 $\pm$ 0.0225 0.4352 $\pm$ 0.0215 0.3831 $\pm$ 0.0658 0.3966 $\pm$ 0.0079
Traffic 0.2007 $\pm$ 0.0014 0.394 $\pm$ 0.0092 0.1339 $\pm$ 0.0166 0.0561 $\pm$ 0.0137 0.5257 $\pm$ 0.0205 0.3833 $\pm$ 0.0271 0.067 $\pm$ 0.0102 0.0653 $\pm$ 0.009
Taxi 0.8948 $\pm$ 0.0017 0.288 $\pm$ 0.0467 0.1756 $\pm$ 0.0056 0.1449 $\pm$ 0.0118 0.5223 $\pm$ 0.0239 0.3106 $\pm$ 0.0512 0.4438 $\pm$ 0.1965 0.5803 $\pm$ 0.2206
Wiki 0.0789 $\pm$ 0.0041 0.1309 $\pm$ 0.0316 0.0649 $\pm$ 0.019 0.0567 $\pm$ 0.0084 0.6588 $\pm$ 0.2225 0.0905 $\pm$ 0.0051 0.142 $\pm$ 0.0624 0.1479 $\pm$ 0.0392
Optimization Training is conducted using the Adam optimizer with a learning rate of $10^{-3}$, and following an ‘LR on plateau’ scheduler (See Pytorch documentation), a weight decay of $10^{-8}$ and $200$ training epochs. Additionally, a separate validation split is used with a temporal size equal to $10$ times the prediction length. When training TimeMCL, the WTA loss (and its variants) was divided by the prediction length $T-t_{0}+1$. All models are trained using gradient norm clipping with a threshold of 10. Training and evaluation are performed in single-precision (float32) with PyTorch.